# The excess of molecular hydrogen in chemical networks without oxygen

### v1

@@ Abstract
pubyear: 2025pagerange: The excess of molecular hydrogen in chemical networks without oxygenData Availability
@@ 1 Introduction
Direct observation of $\mathrm{H}_{2}$ emission is challenging (see Wakelam2017 for a review). Because of its symmetric structure, $\mathrm{H}_{2}$ does not emit or absorb radiation through dipole transitions, which are the most common and easily detectable transitions in molecules. In warm gas ($T\approx 1000\,\mathrm{K}$), collisions can excite $\mathrm{H}_{2}$ molecules to higher vibrational and rotational states. These excited states decay via quadrupole transitions, emitting infrared radiation and enabling radiative cooling. Only if a molecular cloud is heated by a strong radiation field, $\mathrm{H}_{2}$ can be directly observed via fluorescent emission, while it is also in the process of being photo-dissociated (Burkhart2025). Due to the short cooling time at these densities, the majority of $\mathrm{H}_{2}$ is in cold ($\lesssim 100\,\mathrm{K}$) gas. At these temperatures, $\mathrm{H}_{2}$ remains in its ground state and does not emit radiation because quadrupole transitions require higher excitation energies (for $\mathrm{H}_{2}$: $T\geq 511\,\mathrm{K}$).
@@
The contrast between the theoretical simplicity of molecular hydrogen and the observational challenges it presents has led to distinct approaches in studying molecular gas in the Universe. Observational studies rely on indirect tracers, such as CO, to infer the presence and distribution of $\mathrm{H}_{2}$. On the other hand, solving chemical reaction networks with a large number of chemical species is computationally expensive. Theoretical work therefore often uses chemical networks that only contain a small subset of species (e.g. only hydrogen and helium species) and directly model the formation and abundance of $\mathrm{H}_{2}$ without providing information on the CO content. A direct validation of theoretical studies on molecular gas through comparisons with a wide variety of observations is therefore difficult.
@@
In Ploeckinger2025 we introduced Hybrid-Chimes, which is based on the chemical network solver Chimes from chimes2014optthin; chimes2014shielded and implemented in the gravity and hydrodynamics code Swift (swift2024). For computational efficiency, Hybrid-Chimes allows the user to include any selection from metal elements C, N, O, Ne, Si, Mg, S, Ca, and Fe in addition to the primordial elements H and He. While the abundance ratios of the selected elements are evaluated in non-equilibrium and “on-the-fly”, the abundance ratios as well as the cooling and heating rates from the remaining elements are pre-tabulated under the assumption of steady-state chemistry and ionization equilibrium [^1] We use “chemical equilibrium” as short form to steady-state chemistry and ionization equilibrium throughout this work.. For example, in the flagship simulations from the Colibre project (Schaye2025colibre; Chaikin2025a), ten hydrogen and helium species ($e^{-}$, $\mathrm{H}$, $\mathrm{H}^{+}$, $\mathrm{H}^{-}$, $\mathrm{He}$, $\mathrm{He}^{+}$, $\mathrm{He}^{++}$, $\mathrm{H}_{2}$, $\mathrm{H}_{2}^{+}$, and $\mathrm{H}_{3}^{+}$) are evolved in non-equilibrium and their cooling and heating rates are added to those pretabulated from the metal elements (corrected for the non-equilibrium electron abundances, see Ploeckinger2025 for details).
@@
For the specific assumptions on the radiation field, cosmic ray rate, shielding column density and dust content in Ploeckinger2025, we found a systematic over-estimation of $\mathrm{H}_{2}$ when using the minimal chemical network (only H and He species), compared to the results from the full chemical network (11 elements, 157 species). We showed that by adding oxygen to the chemical network (H, He, and O species), the $\mathrm{H}_{2}$ fraction from the full chemical network is recovered. In the following, we discuss some examples to demonstrate that reactions with oxygen species are commonly excluded for the calculation of the abundance of $\mathrm{H}_{2}$.
@@
Grackle (Grackle2017) is a widely used chemical network, implemented e.g. in the gravity / hydro-dynamics codes ChaNGa (Changa2015), Cholla (Cholla2015), Enzo (Enzo2014), Gasoline (Gasoline2017), Gizmo (Gizmo2015), Ramses (Teyssier2002), Arepo (Arepo2020) and SWIFT (swift2024), which solves the non-equilibrium evolution of primordial elements (hydrogen and helium species). In order to account for cooling and heating processes from metal species, especially metal-line cooling, the contributions of metals are calculated and pre-tabulated with Cloudy (last described in cloudy2025), assuming chemical equilibrium. Examples for this approach include the AGORA galaxy simulation comparison project (Agora2014), simulations of isolated galaxies (Emerick2018; Emerick2019; Smith2021), cosmological zoom-in simulations of dwarf galaxies (Revaz2018), and patches of the interstellar medium (Gentry2019). In the cosmological periodic volume simulation Simba-EOR (SimbaEOR2024), the molecular hydrogen fraction is also calculated by Grackle’s primordial chemistry network [^2] In earlier projects (Mufasa: Dave2016 and Simba: Dave2019), the $\mathrm{H}_{2}$ fractions are taken from a subgrid model based on Krumholz2009.. In all these examples, the formation and destruction of molecular hydrogen does not include reactions with gas-phase metals.
@@
The Silcc project (Walch2015) uses a chemical network that follows the evolution of seven chemical species: $e^{-}$, $\mathrm{H}$, $\mathrm{H}^{+}$, $\mathrm{H}_{2}$, $\mathrm{C}^{+}$, $\mathrm{O}$, and $\mathrm{CO}$ and is implemented in the adaptive mesh refinement code Flash (Fryxell2000) to model the evolution of the interstellar medium. While the evolution of some metal species are included in the chemical network, the reactions that form and destroy molecular hydrogen (full list in Micic2012) do not include reactions with metal species. The reactions listed in Micic2012 are also used in Hu2016; Hu2017Grackle; Hu2021 and the Griffin simulations (Lahen2020), again missing the reactions between $\mathrm{H}_{2}$ and metal species. The chemical network Hyacinth (Hyacinth2024) follows selected H, He, C, and O species. Here, $\mathrm{H}_{2}$ can be destroyed by reactions with $\mathrm{C}^{+}$, but not through reactions with oxygen species.
@@
The simulation project Ramses-RTZ (Katz2022) follows the non-equilibrium species abundances of 8 metal elements (C, N, O, Mg, Si, Si, Fe, and Ne) in addition to H and He, coupled to the radiation hydrodynamics solver present in Ramses-RT (Teyssier2002; Rosdahl2015). While they follow the evolution of $\mathrm{H}_{2}$, only reactions of hydrogen and helium species with $\mathrm{H}_{2}$ are included, as described in Katz2017. They benchmark Ramses-RTZ against the results from Cloudy in Katz2022 but exclude molecules for this comparison.
@@
Thesan-Zoom (Kannan2025) follows individual haloes until $z=3$ with the non-equilibrium chemical network from Kannan2020 which evolves the abundances of hydrogen and helium species. They include radiative transfer, particularly of Lyman-Werner radiation, and study e.g. the formation of Pop III stars (Zier2025) and the transition to the formation of Pop II stars. For these low metallicities, metal-line cooling is reduced and cooling from $\mathrm{H}_{2}$ is the dominant cooling process for gas with temperatures of $T\lesssim 10^{4}\,\mathrm{K}$. The destruction of $\mathrm{H}_{2}$ is modeled with collisions of $\mathrm{H}_{2}$ molecular with atomic hydrogen and other hydrogen molecules, as well as photo-ionization and photo-dissociation of $\mathrm{H}_{2}$, but does not include reactions with metal species.
@@
In this paper, we investigate if the $\mathrm{H}_{2}$ excess found in Ploeckinger2025 is limited to the specific model assumptions (radiation field, shielding column density, cosmic ray rate, gas metallicity) and chemical network (Chimes) or if this may indicate a general, systematic overestimation of $\mathrm{H}_{2}$ in reduced chemical networks that include a subset of species and in particular do not include reactions of $\mathrm{H}_{2}$ and oxygen species. We present the method in section 2 and demonstrate the ubiquity of the $\mathrm{H}_{2}$ excess in section 3. In this section (section 3.1) we also show that the $\mathrm{H}_{2}$ excess is drastically reduced when including reactions with oxygen. We discuss implications for studies of both low and high metallicity gas in sections 3.2 and 3.3 and provide a brief summary in section 4. Throughout the paper $\log$ refers to $\log_{\mathrm{10}}$.
@@ 2 Method
We use Cloudy version 25.00, last described in cloudy2025, to calculate the fraction of hydrogen in molecular hydrogen, $f_{\mathrm{H2}}\equiv 2\,n_{\mathrm{H2}}/n_{\mathrm{H}}$, with the molecular hydrogen number density, $n_{\mathrm{H2}}$, and the total hydrogen number density, $n_{\mathrm{H}}$.
For each Cloudy calculation, we assumes that a slab of gas with constant density, temperature, and composition (i.e. element abundances and dust grain properties) is illuminated on one side in a plane-parallel geometry. Cloudy divides the gas into zones with an adaptive thickness. Each subsequent zone is more shielded from the radiation field, as the depth into the cloud, and therefore the shielding column density, $N_{\mathrm{sh}}$, increases. We stop the calculations at a total hydrogen column density of $\log N_{\mathrm{sh,max}}\,[\mathrm{cm}^{-2}]=22.5$ and record the molecular hydrogen fraction, $f_{\mathrm{H2}}$ for each zone.
@@
For the full chemical network, Full, we use the solar abundance ratios from Asplund2009 for all 30 elements included in Cloudy (see Table 1) and scale the total abundance of elements heavier than helium with the relative metallicity $Z/Z_{\odot}$, with $Z_{\odot}=0.0134$ (Asplund2009). Dust grains (silicates and graphites) and PAHs are included with a size distribution that reproduces the overall extinction properties observed in the ISM of the Milky Way Galaxy. This dust model has a dust-to-gas ratio of $\mathcal{DTG}=6.6\times 10^{-3}$ for solar metallicity and is the default dust model in Cloudy. We assume a constant dust-to-metal ratio. Metals are depleted from the gas phase following the strong depletion model ($F_{\star}=1$) from Jenkins2009. While Cloudy does not check for consistency between the grain model and the gas-phase depletion model, we have shown in section 2.1.5 in Ploeckinger2025 that the $F_{\star}=1$ depletion model from Jenkins2009 is consistent with the dust to gas ratio from Cloudy’s ISM dust grain model.
The radiation field illuminating the gas and dust mixture consists of the cosmic microwave background at $z=0$ and an interstellar radiation field (ISRF) whose shape and intensity (Black1987) is representative of the solar neighborhood, and combines the solar vicinity radiation field (Mathis1983) with a Galactic soft X-ray background (Bregman1986). For reference, the strength of this ISRF is $n_{\gamma,\mathrm{ref}}=3.7\times 10^{-4}\,\mathrm{photons\,cm}^{-3}$, where $n_{\gamma,\mathrm{ref}}$ is the number density of hydrogen ionizing photons ($13.6\,\mathrm{eV}$), or $u_{\mathrm{ref}}=1.6\times 10^{-3}\,\mathrm{erg\,cm}^{-2}\,\mathrm{s}^{-1}$, where $u_{\mathrm{ref}}$ is the integrated energy flux in the energy range between 6 and 13.6 eV ($G_{0}=1$ in units of Habing1968).
Finally, a cosmic ray (CR) background is added with a mean hydrogen CR ionization rate of $\zeta_{\mathrm{ref}}=2\times 10^{-16}\,\mathrm{s}^{-1}$ (Indriolo2015) and an $\mathrm{H}_{2}$ secondary ionization rate of $4.6\times 10^{-16}\,\mathrm{s}^{-1}$, following Glassgold1973. We use the converged Cloudy results and the large $\mathrm{H}_{2}$ model from Shaw2005.
@@ 2.1 Model grid
We run Cloudy grids in gas temperature, $T\,[\mathrm{K}]$, and gas density, $n_{\mathrm{H}}$, with different radiation field field strengths ($I_{\mathrm{ISRF}}=n_{\gamma}/n_{\gamma,\mathrm{ref}}=u/u_{\mathrm{ref}}$), cosmic ray ionization rates ($I_{\mathrm{CR}}=\zeta_{\mathrm{CR}}/\zeta_{\mathrm{ref}}$), and metallicities ($Z/Z_{\odot}$), each defined relative to their respective reference values. As both cosmic rays and the interstellar radiation field are related to star formation processes, we assume $I_{\mathrm{ISRF}}=I_{\mathrm{CR}}$. We evaluate each Cloudy calculation at various shielding column densities, $N_{\mathrm{sh}}$, for which we use the values of the zones closest to $N_{\mathrm{sh}}$. Table 2 summarizes the parameter range and variations. The resulting Cloudy input parameter file for each individual run is [^3] We use Cloudy’s grid commands for efficient parallelization but do not include this here for simplicity. We provide the complete scripts to set up and analyze the Cloudy calculations, see the Data Availability Statement for more information.:
@@
constant density
constant temperature [logT] log
hden [lognH] log
stop column 22.5 log
abundances "solar_GASS10.abn" no grains
grains ism [logZ/Zsun]log
grains pah [logZ/Zsun] log
metals [logZ/Zsun] log
metals deplete Jenkins2009 Fstar=1.0 limit 1
CMB redshift 0.0
table ISM [logIISRF] log
cosmic rays background [logICR] log
iterate to convergence
stop temperature off
Database H2
with the grid parameter values from Table 2 for temperature, [logT], density, [lognH], metallicity, [logZ/Zsun], ISRF strength, [logIISRF], and cosmic ray rate, [logICR].

### v2

# arXiv:2512.13447v2[astro-ph.GA] 02 Mar 2026

@@ Abstract
pubyear: 2026pagerange: The excess of molecular hydrogen in chemical networks without oxygenThe excess of molecular hydrogen in chemical networks without oxygen
@@ 1 Introduction
Direct observation of $\mathrm{H}_{2}$ emission is challenging (see Wakelam et al., 2017 for a review). Because of its symmetric structure, $\mathrm{H}_{2}$ does not emit or absorb radiation through dipole transitions, which are the most common and easily detectable transitions in molecules. In warm gas ($T\approx 1000\,\mathrm{K}$), collisions can excite $\mathrm{H}_{2}$ molecules to higher vibrational and rotational states. These excited states decay via quadrupole transitions, emitting infrared radiation and enabling radiative cooling. Only if a molecular cloud is heated by a strong radiation field, $\mathrm{H}_{2}$ can be directly observed via fluorescent emission, while it is also in the process of being photo-dissociated (Burkhart et al., 2025). Due to the short cooling time at these densities, the majority of $\mathrm{H}_{2}$ is in cold ($\lesssim 100\,\mathrm{K}$) gas. At these temperatures, $\mathrm{H}_{2}$ remains in its ground state and does not emit radiation because quadrupole transitions require higher excitation energies (for $\mathrm{H}_{2}$: $T\geq 511\,\mathrm{K}$).
@@
The contrast between the theoretical simplicity of molecular hydrogen and the observational challenges it presents has led to distinct approaches in studying molecular gas in the Universe. Observational studies rely on indirect tracers, such as CO, to infer the presence and distribution of $\mathrm{H}_{2}$. On the other hand, theoretical work often uses chemical networks that only contain a small subset of species (e.g. only hydrogen and helium species) and directly model the formation and abundance of $\mathrm{H}_{2}$ without providing information on the CO content for computational efficiency. A direct validation of theoretical studies on molecular gas through comparisons with a wide variety of observations is therefore difficult.
@@
In Ploeckinger et al. (2025) we introduced Hybrid-Chimes, which is based on the chemical network solver Chimes from Richings et al. (2014a, b) and implemented in the gravity and hydrodynamics code Swift (Schaller et al., 2024). For computational efficiency, Hybrid-Chimes allows the user to include any selection from metal elements C, N, O, Ne, Si, Mg, S, Ca, and Fe in addition to the primordial elements H and He. While the abundance ratios of the selected elements are evaluated in non-equilibrium and “on-the-fly”, the abundance ratios as well as the cooling and heating rates from the remaining elements are pre-tabulated under the assumption of steady-state chemistry and ionization equilibrium [^1] We use “chemical equilibrium” as short form to steady-state chemistry and ionization equilibrium throughout this work.. For example, in the flagship simulations from the Colibre project (Schaye et al., 2025; Chaikin et al., 2025), ten hydrogen and helium species ($e^{-}$, $\mathrm{H}$, $\mathrm{H}^{+}$, $\mathrm{H}^{-}$, $\mathrm{He}$, $\mathrm{He}^{+}$, $\mathrm{He}^{++}$, $\mathrm{H}_{2}$, $\mathrm{H}_{2}^{+}$, and $\mathrm{H}_{3}^{+}$) are evolved in non-equilibrium and their cooling and heating rates are added to those pretabulated from the metal elements (corrected for the non-equilibrium electron abundances, see Ploeckinger et al., 2025 for details).
@@
For the specific assumptions on the radiation field, cosmic ray rate, shielding column density and dust content in Ploeckinger et al. (2025), we found a systematic over-estimation of $\mathrm{H}_{2}$ when using the minimal chemical network (only H and He species), compared to the results from the full chemical network (11 elements, 157 species). We showed that by adding oxygen to the chemical network (H, He, and O species), the $\mathrm{H}_{2}$ fraction from the full chemical network is recovered. In the following, we discuss some examples to demonstrate that reactions with oxygen species are commonly excluded for the calculation of the abundance of $\mathrm{H}_{2}$.
@@
Grackle (Smith et al., 2017) is a widely used chemical network, implemented e.g. in the gravity / hydro-dynamics codes ChaNGa (Menon et al., 2015), Cholla (Schneider and Robertson, 2015), Enzo (Bryan et al., 2014), Gasoline (Wadsley et al., 2017), Gizmo (Hopkins, 2015), Ramses (Teyssier, 2002), Arepo (Weinberger et al., 2020) and SWIFT (Schaller et al., 2024), which solves the non-equilibrium evolution of primordial elements (hydrogen and helium species). In order to account for cooling and heating processes from metal species, especially metal-line cooling, the contributions of metals are calculated and pre-tabulated with Cloudy (last described in Gunasekera et al., 2025), assuming chemical equilibrium. Examples for this approach include the AGORA galaxy simulation comparison project (Kim et al., 2014), simulations of isolated galaxies (Emerick et al., 2018, 2019; Smith, 2021), cosmological zoom-in simulations of dwarf galaxies (Revaz and Jablonka, 2018), and patches of the interstellar medium (Gentry et al., 2019). In the cosmological periodic volume simulation Simba-EOR (Jones et al., 2024), the molecular hydrogen fraction is also calculated by Grackle’s primordial chemistry network [^2] In earlier projects (Mufasa: Davé et al., 2016 and Simba: Davé et al., 2019), the $\mathrm{H}_{2}$ fractions are taken from a subgrid model based on Krumholz et al. (2009).. In all these examples, the formation and destruction of molecular hydrogen does not include reactions with gas-phase metals.
@@
The Silcc project (Walch et al., 2015) uses a chemical network that follows the evolution of seven chemical species: $e^{-}$, $\mathrm{H}$, $\mathrm{H}^{+}$, $\mathrm{H}_{2}$, $\mathrm{C}^{+}$, $\mathrm{O}$, and $\mathrm{CO}$ and is implemented in the adaptive mesh refinement code Flash (Fryxell et al., 2000) to model the evolution of the interstellar medium. While the evolution of some metal species is included in the chemical network, the reactions that form and destroy molecular hydrogen (full list in Micic et al., 2012) do not include reactions with metal species. The reactions listed in Micic et al. (2012) are also used in Hu et al. (2016, 2017, 2021) and the Griffin simulations (Lahén et al., 2020), again missing the reactions between $\mathrm{H}_{2}$ and metal species. The chemical network Hyacinth (Khatri et al., 2024) follows selected H, He, C, and O species. Here, $\mathrm{H}_{2}$ can be destroyed by reactions with $\mathrm{C}^{+}$, but not through reactions with oxygen species.
@@
The simulation project Ramses-RTZ (Katz, 2022) follows the non-equilibrium species abundances of 8 metal elements (C, N, O, Mg, Si, Si, Fe, and Ne) in addition to H and He, coupled to the radiation hydrodynamics solver present in Ramses-RT (Teyssier, 2002; Rosdahl and Teyssier, 2015). While they follow the evolution of $\mathrm{H}_{2}$, only reactions of hydrogen and helium species with $\mathrm{H}_{2}$ are included, as described in Katz et al. (2017). They benchmark Ramses-RTZ against the results from Cloudy in Katz (2022) but exclude molecules for this comparison.
@@
Thesan-Zoom (Kannan et al., 2025) follows individual haloes until $z=3$ with the non-equilibrium chemical network from Kannan et al. (2020) which evolves the abundances of hydrogen and helium species. They include radiative transfer, particularly of Lyman-Werner radiation, and study e.g. the formation of Pop III stars (Zier et al., 2025) and the transition to the formation of Pop II stars. For these low metallicities, metal-line cooling is reduced and cooling from $\mathrm{H}_{2}$ is the dominant cooling process for gas with temperatures of $T\lesssim 10^{4}\,\mathrm{K}$. The destruction of $\mathrm{H}_{2}$ is modeled with collisions of $\mathrm{H}_{2}$ molecular with atomic hydrogen and other hydrogen molecules, as well as photo-ionization and photo-dissociation of $\mathrm{H}_{2}$, but does not include reactions with metal species.
@@
In this paper, we investigate if the $\mathrm{H}_{2}$ excess found in Ploeckinger et al. (2025) is limited to the specific model assumptions (radiation field, shielding column density, cosmic ray rate, gas metallicity) and chemical network (Chimes) or if this may indicate a general, systematic overestimation of $\mathrm{H}_{2}$ in reduced chemical networks that include a subset of species and in particular do not include reactions of $\mathrm{H}_{2}$ and oxygen species. We present the method in section 2 and demonstrate the ubiquity of the $\mathrm{H}_{2}$ excess in section 3. In this section (section 3.1) we also show that the $\mathrm{H}_{2}$ excess is drastically reduced when including reactions with oxygen. We discuss implications for studies of both low and high metallicity gas in sections 3.2 and 3.3 and provide a brief summary in section 4. Throughout the paper $\log$ refers to $\log_{\mathrm{10}}$.
@@ 2 Method
We use Cloudy version 25.00, last described in Gunasekera et al. (2025), to calculate the fraction of hydrogen in molecular hydrogen, $f_{\mathrm{H2}}\equiv 2\,n_{\mathrm{H2}}/n_{\mathrm{H}}$, with the molecular hydrogen number density, $n_{\mathrm{H2}}$, and the total hydrogen number density, $n_{\mathrm{H}}$.
For each Cloudy calculation, we assume that a slab of gas with constant density, temperature, and composition (i.e. element abundances and dust grain properties) is illuminated on one side in a plane-parallel geometry. Cloudy divides the gas into zones with an adaptive thickness. Each subsequent zone is more shielded from the radiation field, as the depth into the cloud, and therefore the shielding column density, $N_{\mathrm{sh}}$, increases. We stop the calculations at a maximum total hydrogen shielding column density, $N_{\mathrm{sh,max}}$, and record the molecular hydrogen fraction, $f_{\mathrm{H2}}$ for each zone. By default, Cloudy evaluates the full gas column once, but the correct total optical depths may be important for line transfer, so we use the option to iterate the calculations of the full shielding column until the optical depths of the total gas column have converged [^3] In rare occasions, a phase transition (e.g. an ionization or dissociation front) coincides with the end of the gas column, which can lead to large relative changes of the total optical depth for each iteration and Cloudy does not find a converged result. See Cloudy documentation for details.. The $\mathrm{H}_{2}$ chemistry in Cloudy is described in detail in Shaw et al. (2005).
We use $\log N_{\mathrm{sh,max}}\,[\mathrm{cm}^{-2}]=22.5$, which ensures that the typical shielding column densities for the H i-$\mathrm{H}_{2}$ transition in the local Universe (e.g. Milky Way Galaxy: $\log N_{\mathrm{H,HI\rightarrow H_{2}}}\,[\mathrm{cm}^{-2}]\approx 21.4$, Wolfire et al., 2008; Small Magellanic Cloud: $\log N_{\mathrm{H,HI\rightarrow H_{2}}}\,[\mathrm{cm}^{-2}]\approx 22.3$, Bolatto et al., 2011) are included for reference. A larger maximum shielding column density would not affect our analysis on the H i-$\mathrm{H}_{2}$ transition, because hydrogen is typically fully molecular at these column densities.
@@
For the full chemical network, Full, we use the solar abundance ratios from Asplund et al. (2009) for all 30 elements included in Cloudy (see Table 1) and scale the total abundance of elements heavier than helium with the relative metallicity $Z/Z_{\odot}$, with $Z_{\odot}=0.0134$ (Asplund et al., 2009). Dust grains (silicates and graphites) and PAHs are included with a size distribution that reproduces the overall extinction properties observed in the ISM of the Milky Way Galaxy. This dust model has a dust-to-gas ratio of $\mathcal{DTG}=6.6\times 10^{-3}$ for solar metallicity and is the default dust model in Cloudy. We assume a constant dust-to-metal ratio. Metals are depleted from the gas phase following the strong depletion model ($F_{\star}=1$) from Jenkins (2009). While Cloudy does not check for consistency between the grain model and the gas-phase depletion model, we have shown in section 2.1.5 in Ploeckinger et al. (2025) that the $F_{\star}=1$ depletion model from Jenkins (2009) is consistent with the dust to gas ratio from Cloudy’s ISM dust grain model.
The radiation field illuminating the gas and dust mixture consists of the cosmic microwave background at $z=0$ and an interstellar radiation field (ISRF) whose shape and intensity (Black, 1987) is representative of the solar neighborhood, and combines the solar vicinity radiation field (Mathis et al., 1983) with a Galactic soft X-ray background (Bregman and Harrington, 1986). For reference, the strength of this ISRF is $n_{\gamma,\mathrm{ref}}=3.7\times 10^{-4}\,\mathrm{photons\,cm}^{-3}$, where $n_{\gamma,\mathrm{ref}}$ is the number density of hydrogen ionizing photons ($13.6\,\mathrm{eV}$), or $u_{\mathrm{ref}}=1.6\times 10^{-3}\,\mathrm{erg\,cm}^{-2}\,\mathrm{s}^{-1}$, where $u_{\mathrm{ref}}$ is the integrated energy flux in the energy range between 6 and 13.6 eV ($G_{0}=1$ in units of Habing, 1968).
Finally, a cosmic ray (CR) background is added with a mean hydrogen CR ionization rate of $\zeta_{\mathrm{ref}}=2\times 10^{-16}\,\mathrm{s}^{-1}$ (Indriolo et al., 2015) and an $\mathrm{H}_{2}$ secondary ionization rate of $4.6\times 10^{-16}\,\mathrm{s}^{-1}$, following Glassgold and Langer (1973).
@@ 2.1 Model grid
We run Cloudy grids in gas temperature, $T\,[\mathrm{K}]$, and gas density, $n_{\mathrm{H}}$, with different radiation field field strengths ($I_{\mathrm{ISRF}}=n_{\gamma}/n_{\gamma,\mathrm{ref}}=u/u_{\mathrm{ref}}$), cosmic ray ionization rates ($I_{\mathrm{CR}}=\zeta_{\mathrm{CR}}/\zeta_{\mathrm{ref}}$), and metallicities ($Z/Z_{\odot}$), each defined relative to their respective reference values. As both cosmic rays and the interstellar radiation field are related to star formation processes, we assume $I_{\mathrm{ISRF}}=I_{\mathrm{CR}}$. We evaluate each Cloudy calculation at various shielding column densities, $N_{\mathrm{sh}}$, for which we use the values of the zones closest to $N_{\mathrm{sh}}$. Table 2 summarizes the parameter range and variations. For reproducibility, we provide the complete scripts to set up and analyze the Cloudy calculations, see the Data Availability Statement for more information.